OpenStudy (anonymous):

-10-3x+2 my answer:-3x=12 do i suppose to divide and get -4

OpenStudy (anonymous):

Are you trying to find x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if so -3x = 8 x = 8/-3 x = -2.67

OpenStudy (anonymous):

this is so when -10-3x+2 = 0 take add the equation with +10 and -2 on both sides That should give you -3x = 8 (as +10-2=8) and so x = -2.67